OpenStudy (anonymous):

Find the product. 1.(2x)(10x) A:20x B:12x C:20x² *** D:5x 2.-3m(m+8) A:-3m²+24 B:-3m²-24m C:-24m² *** D:-4m+5 3.4d(2d-9) A:8d-36d *** B:-d C:-28d² D:8d²-36d 4.Find the area of the rectangle. 7x+1 ________________ | | | | | | 8x |________________| A:15x+1 B:56x²+8x C:56x+8 D:15x²+9x *** 5.find the product. 5b(6b-10) A:11b-10 B:5b²+6b-10 C:-b-10 D:30b²-50b*** Please correct me if I'm wrong on something.

OpenStudy (16shuston):

number 2, 3, and 4 is wrong.

OpenStudy (16shuston):

on number 2 take -3m(m+8)=-3m^2-24m

OpenStudy (16shuston):

number 3, is 4d*2d=8d^2

OpenStudy (16shuston):

and number 4, is 7*8=56

OpenStudy (anonymous):

Number 2 is b

OpenStudy (anonymous):

thank you so much. @16shuston and @KennediSpencer927